# Convex sets and axiom of choice

### v1

@@ 1.2. Faces of convex sets
Here we quickly review some basic facts in the theory of (extreme) faces of convex sets in finite dimensional $\mathbb{R}$}-vector spaces. We will use these facts to develop a general method to find a maximal convex subset of a given subset of a finite dimensional $\mathbb{R}$}-vector space. Facts mentioned here are standard, or are easily deduced from standard facts. The readers are referred to Soltan [14, Chapter 11] for more reading on the theory of extreme faces.
@@ 1.3. Fragments of $\mathrm{AC}$
    For any set $X$, if $|X\times\omega|=|X|$, then $\mathrm{DC}_{X}$ implies $\mathrm{CC}_{X}$.
@@
(4) Assume $\mathrm{DC}_{X}$ and $|X\times\omega|=|X|$. Then we have $\mathrm{DC}_{X\times\omega}$. Let $\langle A_{n}\mid n<\omega\rangle$} be a sequence of nonempty subsets of $X$. Define a binary relation $R$ on $X\times\omega$ by
${\langle a,n\rangle}R{\langle b,m\rangle}\Leftrightarrow b\in A_{m}\land m=n+1.$
Then by applying $\mathrm{DC}_{X\times\omega}$ to $R$ we essentially have a choice function for $\langle A_{n}\mid n_{0}\leq n<\omega\rangle$} for some $n_{0}<\omega$. Thus by adding finite elements we have a choice function for $\langle A_{n}\mid n<\omega\rangle$}.
@@ 8. Higher dimensions
Now we compare $\mathrm{MCV}(V)$’s with more combinatorial fragments of $\mathrm{AC}$.
@@ 9. Summary
The following diagram indicates implications between $\mathrm{MCV}$, $\mathrm{MCV}(V)$ for some $V$’s and other fragments of $\mathrm{AC}$. In the diagram, $(\kappa)$ for a cardinal $\kappa$ denotes $\mathrm{MCV}(\kappa)$. $(\mathrm{WO})$ denotes the statement that $\mathrm{MCV}(\kappa)$ holds for every well-orderable cardinal $\kappa$. $\mathrm{ULF}_{\omega}$ and $\mathrm{LNM}$ respectively denote the statement that every filter on $\omega$ can be extended to an ultrafilter and the existence of a Lebesgue non-measurable subset of $\mathbb{R}$}.
@@ Acknowledgements
The author would like to thank Daisuke Ikegami, Asaf Karagila and Toshimichi Usuba for their helpful comments.

### v2

@@ 1. Introduction
Convex geometry is a subject of mathematics which has relatively heavy dependence on the axiom of choice ($\mathrm{AC}$). For example, some basic theorems of this subject, like the Hahn-Banach theorems and the Krein-Milman theorems need certain fragments of $\mathrm{AC}$ in their proofs (see Bell-Fremlin [2] for the extent of dependence of these theorems on $\mathrm{AC}$).
In this paper, we investigate another, somewhat more naive, connection between convex sets and $\mathrm{AC}$.
@@ 1.2. Faces of convex sets
Here we quickly review some basic facts in the theory of (extreme) faces of convex sets in finite dimensional $\mathbb{R}$}-vector spaces. We will use these facts to develop a general method to find a maximal convex subset of a given subset of a finite dimensional $\mathbb{R}$}-vector space. Facts mentioned here are standard, or are easily deduced from standard facts. The readers are referred to Soltan [16, Chapter 11] for more reading on the theory of extreme faces.
@@ 1.3. Fragments of $\mathrm{AC}$
    (Andretta-Notaro [1]) For any set $X$, if $|X\times 2|=|X|$, then $\mathrm{DC}_{X}$ implies $\mathrm{CC}_{X}$.
@@
For a proof of (4) see [1].
@@ 8. Higher dimensions
Now we compare $\mathrm{MCV}(V)$’s with combinatorial fragments of $\mathrm{AC}$.
@@
We also have some consequences on the existence of bases for some $\mathbb{R}$}-vector spaces.
Proposition 8.4.
For each $\mathbb{R}$}-vector space $V$, $\mathrm{SC}_{V}$ implies that every linearly independent subset of $V$ can be extended to a basis of $V$.
Proof.
Assume $\mathrm{SC}_{V}$ and let $B_{0}$ be any linearly independent subset of $V$. Let
$C=\{S\in[V\setminus B_{0}]^{<\omega}\mid\text{$S\cup B_{0}$ is linearly independent}\}.$
Then $C$ is a subset of $V]^{<\omega}$ and $\emptyset\in C$ holds. Now apply $\mathrm{SC}_{V}$ to obtain a maximal $P\subseteq V$ such that $P]^{<\omega}\subseteq C$. Then it is easy to see that $P\cup B_{0}$ is a basis of VV.∎
By Lemma 8.2(1) and Proposition 8.4 we have
Corollary 8.5.
For each $\mathbb{R}$}-vector space $V$, if $\mathrm{MCV}(|V|)$ holds then every linearly independent subset of $V$ can be extended to a basis.∎
By the above corollary and the remark below the proof of Proposition 8.1 we have the following.
Corollary 8.6.
Suppose $\mathrm{MCV}(V)$ holds for either $V={\mathbb{R}}_{\mathbb{R}}$, ${\mathbb{R}}^{\omega}$, $l^{\infty}$, $C({\mathbb{R}},{\mathbb{R}})$ or $\mathcal{H}$. Then these spaces are all isomorphic (as $\mathbb{R}$}-vector spaces).∎
@@ 9. Summary
The following diagram indicates implications between $\mathrm{MCV}$, $\mathrm{MCV}(V)$ for some $V$’s and other fragments of $\mathrm{AC}$. In the diagram, $(\kappa)$ for a cardinal $\kappa$ denotes $\mathrm{MCV}(\kappa)$. $(\mathrm{WO})$ denotes the statement that $\mathrm{MCV}(\kappa)$ holds for every well-orderable cardinal $\kappa$. For an $\mathbb{R}$}-vector space $V$, $\mathrm{BS}_{V}$ denotes the statement that every linearly independent subset of $V$ can be extended to a basis of $V$. $\mathrm{ULF}_{\omega}$ and $\mathrm{LNM}$ respectively denote the statement that every filter on $\omega$ can be extended to an ultrafilter and the existence of a Lebesgue non-measurable subset of $\mathbb{R}$}.
@@
Question 9.5.
Does $\mathrm{MCV}(V)$ imply that $V$ has a basis? By Corollary 8.5 $\mathrm{MCV}(|V|)$ does, but it is not clear if $\mathrm{MCV}(V)$ does. Is there any implication of the other direction, that is, is $\mathrm{MCV}(V)$ derived from the existence of a basis for some $\mathbb{R}$}-vector space?
@@ Acknowledgements
The author would like to thank Asaf Karagila, who read an earlier version of this paper and made many useful comments, calling the author’s attention to some papers to reference and to the question how $\mathrm{MCV}$-type axioms relates to the existence of bases for $\mathbb{R}$}-vector spaces. We would like to thank Daisuke Ikegami and Toshimichi Usuba for their helpful comments as well.